Partners in a busness agreed to take out two loans totaling $35000. The yearly interest rates were 12% and 15%, and the total yearly interest was $4650. Find the amount of each loan.
Answer by tutorcecilia(2152)

Ax+By=C [use the standard form of a line]
.12x+.15y=4650 [change percents to decimals]
and x+y=35,000
So,x+(35,000-x)=35,000
Putting it all together:
.12x+.15(35,000-x)=4650 [simplify and solve for x]
.12+5250-.15x=4650
-.03x=4650-5250
-.03x=-600
x=$20,000 invested at 12%
y=$35,000-x=($35,000-$20,000)=$15,000 invested 15%
.
Check by plugging in all of the values back into the original equation.